# Genetic Data Governance in Crisis: Policy Recommendations for Safeguarding Privacy and Preventing Discrimination

### v1

@@ 3.2.2. Genetic Data Privacy Laws
Many states have genetic data privacy laws, either as a standalone law or as part of a broader data privacy law. These laws [^10] Some of these laws have also led to a change in company services in that state. For example, the passing of Illinois’ BIPA (Biometric Information Protection Act) resulted in Amazon and Google removing certain services of Alexa and Google Photos respectively after lawsuits which claimed that the company violated BIPA. were generally passed after GINA with the rise of DTC genetic testing companies. Standalone genetic data privacy laws tend to be written specifically to cover DTC genetic testing companies and are often extensions of genetic non-discrimination laws passed around the time of GINA.
@@
Both data privacy and genetic privacy laws have similar provisions (Institute, 2025), including a customer right to request deletion of genetic data and samples and the necessity of consent for data transfers. Some states allow consumers to bring lawsuits for violations of the privacy law (a private right of action). In general though, privacy laws do not apply to deidentified genetic data (as per HIPAA). An important contrast to the US federal and state bills on privacy is the EU General Data Protection Regulation (GDPR). GDPR not only regulates all personal data within the EU but also the transfer of data outside of the EU, data minimization, and compliance. Genetic data is labeled as “special category data” in GDPR as it has many identifiers associated with it.
4. Risk Assessment Framework
@@
Our goal in this section is to develop a risk assessment framework for genetic data governance. Similar to the aims of the Blueprint for the AI Bill of Rights (Park, 2023), we motivate our framework with three central questions: (1) What values (i.e., moral principles and civil liberties) should be preserved? (2) What are the vulnerabilities in the current system that can compromise these values? (3) What are the specific harms that can result from the vulnerabilities?
4.1. What values should be preserved?
We seek to identify values that are future-proof and can be encoded in a system of genetic data collection. By examining the Four Pillars, six key values emerge:
  1. (1)
@@ 4.1. What values should be preserved?
    Right to action: The individual, and the individual only, has the choice to submit (and the freedom to not submit) their genome [^11] We use “genome” to refer to a physical DNA samples of an individual and any sequencing/genotyping data..
  2. (2)
@@
  5. (5)
    Protecting opportunities for advancement: Genetic data should not be used to deprive the individual of opportunities in any domain, including (but not limited to) education, access to financial tools, health insurance, housing, social services, and reproductive choices.
  6. (6)
    Benefits of inclusion: The Belmont principle states “those who bear the burdens of research (i.e., those who are exposed to the discomforts, inconveniences, and risks) should receive the benefits in equal measure to the burdens”(U.S. Department of Health & Human Services (1979), HHS). We believe that this applies to individuals and their genetic data. [^12] Variant Bio is a biotech company that collects genetic data from Indigenous groups from around the world and participates in revenue sharing with the communities they collect data from.
4.2. What are the vulnerabilities in the current system that can compromise these values?
@@
  1. (1)
    Unsettled science: The role of genetics in shaping complex traits is poorly understood. As a consequence, the deterministic nature of genetics may be overstated, particularly for behavioral and cognitive traits.
  2. (2)
    Rapid evolution of genetic data/methods: As genetic data collection, sequencing, and methodologies quickly evolve, legal protections fall behind. For example, education discrimination is not included in GINA, which was passed five years before the first large GWAS on educational attainment (Rietveld et al., 2013).
  3. (3)
@@
4.3. What are the harms that arise as a result?
Lastly, we will outline several consequences that could arise from violations of our key values through the above vulnerabilities. These downstream effects, which we refer to as harms, can affect an individual and their immediate genetic relatives and can also have a broader effect on communities and the population at large.
  1. (1)
    Leakage to the family: The genetic dragnet means that any conclusion about an individual’s genetic trait can be linked back to their relatives, even if no action is taken by their relatives in either (1) submitting data or (2) opting in to receive information. Any derived secondhand knowledge can immediately impact insurance (health, life, etc.) and medical treatment plans, as well as compromise identity.
  2. (2)
    Loss of anonymity: It is straightforward to identify some traits about an individual from their genetic data, and subsequently from their genetic relatives. Common examples include: race, gender, ethnicity, and markers for certain diseases. Leaky genetic data interfaces can thus compromise information about an individual, their genetic relatives, and unborn children that the individual (and implicated relatives) may wish to maintain as private. As the refinement of reference datasets continues and data collection grows [^13] It has been estimated that, in a sample of three million European-descent Americans, 99% of individuals would have a least one third cousin in the dataset (Erlich et al., 2018)., anonymity becomes even more elusive.
  3. (3)
    Loss of Data Control: Private institutions offering genetic tests control the entire lifecycle of individuals’ data from collection and storage to monetization. This control creates significant transparency issues, as it is often unclear how data is monetized, which parties benefit, or what agreements include an individual’s data. Users may opt out of participation, but unregulated data flows ensure that their genetic footprint persists indefinitely. This lack of transparency is particularly concerning in cases of bankruptcy or acquisition. For example, in 2024, Tempus AI acquired Ambry Genetics, a company that sequences genetic data for 400,000 patients annually (Tempus, 2025). Tempus stated its intent to ”leverage this data and augment Tempus’ current data offering,” raising questions about how patient data is repurposed and monetized in such deals. Ultimately, an individual’s decision to submit genetic data for testing often entails relinquishing full ownership of their genetic information. Without robust safeguards, privacy risks emerge from opaque data practices and leaky interfaces, leaving individuals vulnerable to misuse or unauthorized access to their genetic data.
  4. (4)
@@ 5. Case Studies
We apply our risk assessment framework to four case studies which highlight regulatory gaps across the Four Pillars. These case studies, ranging from past events to speculative scenarios, all involve privacy loss or potential discrimination. We choose case studies that illustrate genetic data leakage in both, a pillar-to-pillar and one-to-many pillar settings.
5.1. Genetics and Education
@@
Polygenic scores (PGS) are popular in the social/behavioral sciences for predicting social outcome traits, such as educational attainment (EA; number of schooling years completed by an adult). We will use EA as the example throughout this case study, but note that there is also interest in predicting standardized testing scores, performance in mathematics, and other traits with substantial environmental influences.
@@
A common metric for assessing PGS accuracy is the percentage of trait variance it explains: higher percentages indicate better predictive performance. A recent 23andMe EA PGS, based on data from over 3 million customers of European descent, explains 12-16% of the variance  [^14] As a useful comparison, mother’s education explains 15% of the variance in EA (Lee et al., 2018) in educational attainment (EA) (Okbay et al., 2022). This effectively means that 50-70% of individuals with PGS scores in the top 10% for EA are predicted to graduate college. However, the PGS accuracy significantly decreases when applied to African American customers. This is an example of the commonly observed “portability problem” (Martin et al., 2017), where a PGS derived from GWAS in one population predicts poorly in another due to confounding [^15] Confounding in GWAS can be genetic (non-causal variants correlated with causal ones) or environmental (non-causal variants correlated with causal environmental factors), leading to potential statistical artifacts in effect sizes..
@@
There have been several calls to use EA PGS to inform education policy. Harden et al. propose the use of math-performance PGS to identify “leaks” in the education system: for example, by identifying high math PGS students who perform poorly, they claim educators could pinpoint why and how students are failing to reach their potential [^16] An example they use: 31% of high PGS students in good schools take calculus, compared to 24% of the same-scoring students in poor-performing schools. (Harden et al., 2020). Plomin & von Stumm take it further: they use the term “precision education” (akin to precision medicine) to propose a tailor-made, individualized education that is genetics-informed (R and S, 2018). Statements like this, combined with statements such as “students with higher polygenic scores for years of education have, on average, higher cognitive ability, better grades and come from families with higher SES [socioeconomic status]” (Smith-Woolley et al., 2018) are cause for concern because they invoke a sense of genetic determinism. However, other predictors (parents’ educational status, socioeconomic status) explain similar amounts of variance in EA (Lee et al., 2018; Morris et al., 2020) and—unlike DNA—are mutable through social policy changes.
@@
Several of our values would be violated if children were required to submit their DNA (Right to Action) or educational opportunities were denied to children based on their genetic potential (Opportunities for Advancement). Through the vulnerabilities of unsettled science and the rapid evolution of genetic methods, harms such as leakage to the family can occur and affect not only children but their families and future.
5.2. DTC Genetic Data Brokerage and Leakage
Several FDA-approved 23andMe tests17171723andMe tests with FDA 510(k) clearance are verified as substantially equivalent to existing devices or methods. Only health-related tests, such as those identifying genetic risks, carrier status, or drug responses, qualify for this review. provide customers with results containing sensitive information, including carrier screening reports, genetic health risk assessments, and BRCA1/BRCA2 variant analyses. In addition, they offer genealogy (to identify relatives) and ancestry services. In an attempt to protect customer privacy, 23andMe offers an opt-in/opt-out policy when it comes to sharing their samples in studies with third-parties.
In 2023, hackers accessed 14,000+ 23andMe user accounts by using recycled login information (credential stuffing). Due to the relatives linkage feature in genetic data reports, the hackers were able to then access information about an additional 5.5M users (23a, 2023b); there were also reports that the hacker targeted Ashkenazi Jewish and Chinese ancestry profiles (Carballo et al., 2024). Among the many values violated, the Right of Privacy and Ownership of the Genome are particularly relevant to this case as 23andMe did not immediately notify compromised users or the public. Several class action lawsuits were filed (e.g., (for the Northern District of California, 2023)), with plaintiffs complaining that 23andMe failed to adequately protect their sensitive information. 23andMe responded by enforcing multifactor authentication (MFA) of user accounts, which was previously optional (Whittaker, 2023), and a serious technological vulnerability. A key gap that allowed for this issue to occur is the lack of consensus around security responsibility, which resulted in the disproportionate and harmful targeting of underrepresented groups. The 23andMe breach highlights the consequences of the relational nature of genetic data, where an individual’s information can be compromised through poor public protections, even if they personally take precautions. As one Reddit user quipped “Your genetic data is only as secure as your relatives’ passwords.”
5.3. Genetic Data Collection of Detained Noncitizens
@@
A recent Georgetown University report from the Center of Privacy and Technology (CPT) examines the U.S. federal government’s practice of collecting DNA from detained migrants (Glaberson et al., 2024). This practice, which began with the 2005 DNA Fingerprint Act and expanded significantly in 2020—a mandate that the Department of Homeland Security must collect DNA from all detainees, even briefly detained [^18] States have varying rules about DNA collection from suspects and detainees: 28 states allow DNA to be taken from someone before they are convicted of any crime and most states restrict DNA collection to cases involving felony charges or convictions, rather than allowing collection for any arrest or detention (Samuels et al., 2012).—has led to a dramatic increase in detainee representation in CODIS (the federal DNA database), rising from 0.21% in 2019 to 9.21% in 2023. By 2020, approximately 25,000 noncitizens were added to the database under the “detainee” classification.
@@
The CPT report highlights several concerning aspects of this practice, particularly its disproportionate impact on migrants of color and issues of consent. In our Risk Assessment Framework, this immediately violates the Right to Action and Right to Knowledge. Many migrants undergo DNA collection without understanding its implications, sometimes believing it to be a COVID-19 test or submitting under threat of criminal prosecution. The report argues that this practice violates the Fourth Amendment in collecting DNA from detainees without probable cause. Additionally, even though CODIS was designed to be privacy-protective by collecting only 20 markers—a tiny slice of the genome believed to be medically neutral—this limited data can still identify relatives through partial matching and, as an analysis 25 years later would show, are slightly informative of disease risk (Bañuelos et al., 2022). The expansive collection of genetic data has far-reaching implications through the guilt-by-association vulnerability. Since CODIS records are difficult to expunge, they can restrict Opportunities for Advancement not only for detainees but also for their relatives and descendants.
5.4. Underwriting Life Insurance with AI and Genetic Data
@@
Life insurance underwriting is the process whereby an insurance company uses personal and health information to assess the risk of insuring an applicant. The relationship between applicants and insurance companies is already fraught: (Devnos et al., 2016) find that patients are more likely to share their genetic info with Google than insurance companies. The future of life insurance underwriting is expected to become increasingly computational and automated via the usage of AI and the collection of more personalized, individual genetic data (Filabi and Duffy, 2021; Balasubramanian et al., 2021; Rothstein, 2018; Koleva-Kolarova et al., 2022). While medical records and demographic factors can be used for mortality analysis, the inclusion of genetic factors can mean that risk prediction can be performed much earlier in an applicant’s life without necessarily the same amount of records as an older individual, emphasizing the risk of genetic discrimination (Karlsson Linnér and Koellinger, 2022) based on the potential to have risk factors (e.g., a genetic risk for high blood pressure versus clinically-measured high blood pressure). Life insurance is not covered by GINA and life insurance companies can access medical records (which may include genetic test results) as part of an application. Bills have been introduced in several states that would restrict the use of genetic information in underwriting, but these efforts largely failed: As of 2022, out of thirty-seven proposed bills across all states, three were introduced, eight were signed by the governor, demonstrating the geographical dependency of policy-based protection (Rickard, 2022).
@@
Life insurance companies may be interested in using polygenic scores (PGS) to assess an individual’s risk for various diseases. Indeed, in early 2024, U.S.-based life insurance company MassMutual and U.K.-based Genomics plc announced a partnership, offering free genetic testing to MassMutual’s life insurance customers (MassMutual, 2024). However, the press release stressed that MassMutual would not receive individual results and that current premiums/policies would be unaffected (MassMutual, 2024). This example particularly highlights the leaky interface between pillars, where the Right of Privacy and Right of Action are called into question. AI insurance underwriting algorithms already suffer from racial biases (Lee, 2022). That PGS suffer from the portability problem (itself caused by systematic biases in training datasets) means that these biases could be perpetuated in genetics-informed life insurance underwriting. Additionally, given the relational aspect of genetic data, genetic-based underwriting could affect other biological relatives’ applications, harming not only relatives but any others with the same genetic markers that are associated with mortality risks. The potential harms of using genetic data as a part of underwriting in life insurance are one such example of violations of genetic privacy with the rapid evolution and usages of genetic data with unsettled science.
6. Recommendations
We argue that for public infrastructure and private companies to productively and ethically make use of genetic data, several amendments need to be made to the current genetic data governance system. Our three recommendations address open privacy concerns, legislative scoping for policy changes, and best practices for bodies handling genetic data.
6.1. Recommendation 1: Redefining Genetic Data
Issue: Legal policy surrounding genetic privacy notably excludes deidentified or anonymized data from protection.
Recommendation: Given that we argue that genetic data is unique compared to any other identifying data, we suggest genetic data be defined using the following language:
@@ 6.1. Recommendation 1: Redefining Genetic Data
“Genetic data refers to any information relating to an individual’s genetic characteristics, including but not limited to DNA or RNA sequences, gene expression profiles, or any data derived from a biological sample—including that of a relative—regardless of format. This data is always considered identifiable by its nature—or personally identifiable information (PII)—as it pertains to unique biological attributes that can potentially be linked to a specific individual, their biological relatives, or identifiable group. For the purposes of this definition, any de-identified, pseudonymized, or anonymized genetic information shall be treated as genetic data, regardless of measures taken to mask individual identities, recognizing that the inherent characteristics of genetic information may enable re-identification through advanced technological or data cross-referencing methods.”
6.2. Recommendation 2: Extending Protections for Genetic Discrimination
@@
Issue: GINA is a vital piece of federal legislation that protects against genetic discrimination in employment and health insurance domains. However, other domains in which there exists potential for genetic discrimination are not protected by GINA: other insurance domains (life, long-term care, disability), housing, and education.
@@
Recommendation: We recommend that additional federal laws should be enacted to extend GINA’s coverage beyond employment and health insurance. The most comprehensive state genetic anti-discrimination law is California’s CalGINA, which includes housing, mortgage brokerage, education, and more. However, despite these extensions, CalGINA does not cover life, disability, or long-term care insurance. We suggest that extension bills for GINA should explicitly cover life, disability, long-term care, as well as education and any other opportunities for advancement to be protected against genetic discrimination. It should be written in such a way that prohibits any barrier to opportunity, even those not anticipated at the time of writing. Additionally, legislation should be written to explicitly bar genetic risk for complex traits known to have significant environmental influences (such as cardiovascular disease) from being considered a preexisting condition.
6.3. Recommendation 3: A Genetic Data Regulation Framework
Issue: Current regulations were designed to govern one application of genetic data—or Pillar—at a time. This has led to “leaky protections”, where the use of genetic data in one Pillar can affect opportunities and decisions in other Pillars (e.g., clinical tests being used in life insurance).

### v2

@@ Introduction
Many states have genetic data privacy laws, either as a standalone law or as part of a broader data privacy law. These laws were generally passed after GINA with the rise of DTC genetic testing companies. Standalone genetic data privacy laws tend to be written specifically to cover DTC genetic testing companies and are often extensions of genetic non-discrimination laws passed around the time of GINA.
@@
Both data privacy and genetic privacy laws have similar provisions (Prince et al. 2021), including a customer right to request deletion of genetic data and samples and the necessity of consent for data transfers. Some states allow consumers to bring lawsuits for violations of the privacy law (a private right of action). In general though, privacy laws do not apply to deidentified genetic data (as per HIPAA). An important contrast to the US federal and state bills on privacy is the EU General Data Protection Regulation (GDPR). GDPR not only regulates all personal data within the EU but also the transfer of data outside of the EU, data minimization, and compliance. Genetic data is labeled as “special category data” in GDPR as it has many identifiers associated with it.
Blindspots for Regulatory Bodies Governing Genetic Data Collection and Usage
A critical tension within the current protective framework, predating even advanced AI applications, lies between traditional single-gene tests and newer probabilistic measures of genetic risk, notably polygenic scores (PGS). Legislation like GINA was largely envisioned for single-gene tests that detect specific variants directly linked to conditions (e.g., BRCA1/2 for breast cancer, HTT for Huntington’s). These tests offer relatively deterministic insights, which existing laws sought to shield from discriminatory use.
However, PGS operate differently. They don’t “detect” individual genotypes or mutations in the direct sense that GINA outlines in its definition of “genetic information” (an analysis that “detects genotypes, mutations, or chromosomal changes” (Congress 2008)). Instead, PGS represent a secondary computational analysis, transforming raw genetic data into statistical models of relative risk across populations. This transformation from direct detection to probabilistic, interpretive analysis arguably places PGS outside GINA’s explicit protections for employment and health insurance decisions (Callier and Prince 2024; Yanes et al. 2024). This potential loophole is significant because the law focuses on the detection of genetic information rather than the interpretive analyses and risk scores derived from it.
This distinction also manifests in regulatory oversight. The FDA, for instance, regulates well-defined single-gene tests with clear clinical actions, but PGS largely fall outside its purview. While the FTC might address PGS claims under fraud or deceptive practice statutes, this offers a different, and potentially less direct, form of consumer protection than the specific anti-discrimination mandates of GINA or the clinical validation oversight of the FDA.
Risk Assessment Framework
@@
Our goal in this section is to develop a risk assessment framework for genetic data governance. We adopt the structure of the aims of the Blueprint for the AI Bill of Rights (Park 2023) and motivate our framework with three central questions: (1) What values (i.e., moral principles and civil liberties) should be preserved? (2) What are the vulnerabilities in the current system that can compromise these values? (3) What are the specific harms that can result from the vulnerabilities?
What values should be preserved?
We seek to identify values that are future-proof and can be encoded in a system of genetic data collection, considering practical ideas of control, consent, and non-discrimination surrounding genetic data. Using Viljoen’s framework of the relational theory of data governance (Viljoen 2024), we identify vertical relationships (between data collectors, i.e. the Four Pillars, and individuals) and horizontal relationships (between genetically connected individuals). This framework must protect personal and civil liberties for both data submitters and their implicated genetic relatives, encompassing both vertical and horizontal relationality. Through this analysis, six key values emerge. The first four values are drawn particularly from Nissenbaum’s framework of contextual integrity, highlighting genetic data transfer while preserving personal freedoms (Nissenbaum 2019, 2004; Barth et al. 2006).
  1. 1.
@@
    Right to action: The individual, and the individual only, has the choice to submit (and the freedom to not submit) their genome.
  2. 2.
@@
  5. 5.
    Protecting opportunities for advancement: Genetic data should not be used to deprive the individual of opportunities in any domain, including (but not limited to) education, access to financial tools, health insurance, housing, social services, and reproductive choices, thus preventing a “genetic underclass” (Ajunwa 2016).
  6. 6.
    Benefits of inclusion: The Belmont principle states “those who bear the burdens of research (i.e., those who are exposed to the discomforts, inconveniences, and risks) should receive the benefits in equal measure to the burdens” (U.S. Department of Health & Human Services (1979) HHS). We believe that this applies to individuals and their genetic data.
What are the vulnerabilities in the current system that can compromise these values?
@@
  1. 1.
    Unsettled science: The role of genetics in shaping complex traits is poorly understood, and integrating genetic data into multi-modal AI models will likely not improve this understanding. As a consequence, the deterministic nature of genetics may be overstated, particularly for behavioral and cognitive traits.
  2. 2.
    Rapid evolution of genetic data/methods: As genetic data collection, sequencing, and methodologies quickly evolve, legal protections fall behind. For example, education discrimination is not included in GINA, which was passed five years before the first large GWAS on educational attainment (Rietveld et al. 2013).
  3. 3.
@@
What are the harms that arise as a result?
In the final section of our risk assessment framework, we outline the consequences that could arise from violations of our key values through the above vulnerabilities. These downstream effects, which we refer to as harms, can not only affect an individual and their immediate genetic relatives and but can also have a broader effect on communities and the population at large.
  1. 1.
    Leakage to the family: Any conclusion about an individual’s genetic trait may be linked back to their relatives, even if no action is taken by their relatives in either (1) submitting data or (2) opting in to receive information. Any derived secondhand knowledge can immediately impact insurance (health, life, etc.) and medical treatment plans, as well as compromise identity.
  2. 2.
    Loss of anonymity: It is straightforward to identify some traits about an individual from their genetic data, and subsequently from their genetic relatives. Common examples include: race, gender, ethnicity, and markers for certain diseases. As the refinement of reference datasets continues and data collection grows, leaky genetic data interfaces can thus compromise information about an individual, their genetic relatives, and unborn children that the individual (and implicated relatives) may wish to maintain as private.
  3. 3.
    Loss of Data Control: Private institutions offering genetic tests control the entire lifecycle of individuals’ data from collection and storage to monetization. This control creates significant transparency issues, as it is unclear how data is monetized, which parties benefit, or what agreements include an individual’s data. Users may opt out of participation, but unregulated data flows ensure that their genetic footprint persists indefinitely. This lack of transparency is particularly concerning in cases of bankruptcy or acquisition. For example, in 2024, Tempus AI acquired Ambry Genetics, a company that sequences genetic data for 400,000 patients annually (Tempus 2025). Tempus stated its intent to “leverage this data and augment Tempus’ current data offering,” raising questions about how patient data is repurposed and monetized in such deals. Ultimately, an individual’s decision to submit genetic data for testing often entails relinquishing full ownership of their genetic information. Without robust safeguards, privacy risks emerge from opaque data practices and leaky interfaces, leaving individuals vulnerable to misuse or unauthorized access.
  4. 4.
@@
We apply our risk assessment framework to four case studies which highlight regulatory gaps across the Four Pillars. These case studies, ranging from past events to speculative scenarios, all involve privacy loss or potential discrimination. We explore case studies that illustrate genetic data leakage in both, a pillar-to-pillar and one-to-many pillar settings.
Genetics and Education
@@
Polygenic scores (PGS) are popular in the social/behavioral sciences for predicting social outcome traits, such as educational attainment (EA; number of schooling years completed by an adult). We will use EA as the example throughout this case study, but note that there is interest in predicting standardized testing scores, mathematics performance, and other traits with substantial environmental influences.
@@
A common metric for assessing PGS accuracy is the percentage of trait variance it explains: higher percentages indicate better predictive performance. A recent 23andMe EA PGS, based on data from over 3 million customers of European descent, explains 12-16% of the variance in educational attainment (EA) (Okbay et al. 2022). This effectively means that 50-70% of individuals with PGS scores in the top 10% for EA are predicted to graduate college. However, the PGS accuracy significantly decreases when applied to African American customers. This is an example of the commonly observed “portability problem” (Martin et al. 2017), where a PGS derived from GWAS in one population predicts poorly in another due to confounding.
@@
There have been several calls to use EA PGS to inform education policy. Harden et al. propose the use of math-performance PGS to identify “leaks” in the education system: for example, by identifying high math PGS students who perform poorly, they claim educators could pinpoint why and how students are failing to reach their potential (Harden et al. 2020). Plomin & von Stumm take it further: they use the term “precision education” (akin to precision medicine) to propose a tailor-made, individualized education that is genetics-informed (Plomin and von Stumm 2018). Statements like this, combined with statements such as “students with higher polygenic scores for years of education have, on average, higher cognitive ability, better grades and come from families with higher SES [socioeconomic status]” (Smith-Woolley et al. 2018) are cause for concern because they invoke a sense of genetic determinism. However, other predictors (parents’ educational status, socioeconomic status) explain similar amounts of variance in EA (Morris, Davies, and Davey Smith 2020; Mostafavi et al. 2020) and—unlike DNA—are mutable through social policy changes.
@@
Several of our values would be violated if children were required to submit their DNA (Right to Action) or educational opportunities were denied to children based on their genetic potential (Opportunities for Advancement). Through the vulnerabilities of unsettled science and the rapid evolution of genetic methods, leakage to the family can occur and affect not only children, but their families and future. Crucially, these profound challenges exist even before considering the added complexities and potential for misuse introduced by integrating such scores with advanced AI systems.
DTC Genetic Data Brokerage and Leakage
The fragility of genetic data governance is starkly illustrated by direct-to-consumer (DTC) company 23andMe. Beyond FDA-approved genetic tests (e.g., BRCA1/2 variants), carrier screenings, health risk assessments, and genealogy services, the company also collects extensive non-genetic survey data and conducts its own GWAS.
In 2023, a credential stuffing attack compromised 14,000+ user accounts. Exploiting 23andMe’s relative linkage feature hackers accessed data for an additional 5.5M users (23andMe 2023b) in a powerful example of genetic data’s relationality. The hackers also reportedly targeted Chinese and Ashkenazi Jewish profiles (Carballo, Schmall, and Tumin 2024). A critical factor enabling this breach was the lack of consensus around security responsibility, which resulted in the disproportionate and harmful targeting of underrepresented groups. The breach highlighted how an individual’s genetic privacy is hostage to others’ security practices, aptly summarized by a Reddit user: “Your genetic data is only as secure as your relatives’ passwords.”
The security and integrity of consumer genetic data was once again compromised when 23andMe filed for bankruptcy and announced its database and assets were for sale (23andMe 2025). This prompted widespread calls for data deletion (Eddy 2025) and a federal probe into the sale’s privacy and national security risks (U.S. House Committee on Oversight and Government Reform 2025). The situation culminated on May 19th, 2025, with the announcement that pharmaceutical giant Regeneron had acquired 23andMe’s assets for \$256M (Regeneron Pharmaceuticals, Inc. 2025). Regeneron thus gained access to existing genetic data, extensive non-genetic information, model weights, and stored biological samples, which allow for future whole-genome sequencing. Regeneron, in compliance with the terms of the sale (23andMe, Inc. 2025), has pledged to uphold 23andMe’s terms and conditions (T&Cs), which include the right to delete data (Regeneron Pharmaceuticals, Inc. 2025).
The 23andMe saga reveals critical vulnerabilities and rights violations within our risk assessment framework. The data breach infringed on Right to Privacy, while the asset sale to Regeneron highlighted contested Ownership of the Genome, as users’ deletion rights via T&Cs contrast with the de facto transfer of dataset ownership. This underscores the dangers of a fragmented regulatory landscape: Californians have protections during third-party sales, but those in states like Pennsylvania lack equivalent federal or state safeguards for DTC genetic data sales. Finally, Regeneron’s acquisition of biological samples and non-genetic data, especially for use in new AI models, exposes vulnerabilities from rapidly evolving technology and unsettled science, enabling applications far beyond users’ original consent or understanding.
Genetic Data Collection of Detained Noncitizens
@@
A recent Georgetown University report from the Center of Privacy and Technology (CPT) examines the U.S. federal government’s practice of collecting DNA from detained migrants (Glaberson, Tse, and Tucker 2024). This practice, which began with the 2005 DNA Fingerprint Act and expanded significantly in 2020—a mandate that the Department of Homeland Security must collect DNA from all detainees, even briefly detained—has led to a dramatic increase in detainee representation in CODIS (the federal DNA database), rising from 0.21% in 2019 to 9.21% in 2023. By 2020, approximately 25,000 noncitizens were added to the database under the “detainee” classification.
@@
The CPT report highlights several concerning aspects of this practice, particularly its disproportionate impact on migrants of color and issues of consent. In our Risk Assessment Framework, this immediately violates the Right to Action and Right to Knowledge. Many migrants undergo DNA collection without understanding its implications, sometimes believing it to be a COVID-19 test or submitting under threat of criminal prosecution. The report argues that this practice violates the Fourth Amendment in collecting DNA from detainees without probable cause. Additionally, even though CODIS was designed to be privacy-protective by collecting only 20 markers—a tiny slice of the genome believed to be medically neutral—this limited data can still identify relatives through partial matching and, as an analysis 25 years later would show, are slightly informative of disease risk (Bañuelos et al. 2022). The expansive collection of genetic data has far-reaching implications through the guilt-by-association vulnerability. Since CODIS records are difficult to expunge, they can restrict Opportunities for Advancement not only for detainees but also for their relatives and descendants.
Underwriting Life Insurance with AI and Genetic Data
@@
Life insurance underwriting is the process whereby an insurance company uses personal and health information to assess the risk of insuring an applicant. The relationship between applicants and insurance companies is already fraught: (Devnos, Wang, and Iyer 2016) found that patients are more likely to share their genetic information with Google than insurance companies. The future of life insurance underwriting is expected to become increasingly computational and automated through the usage of AI and the collection of more personalized, individual genetic data (Balasubramanian, Libarikian, and McElhaney 2021; Filabi and Duffy 2021; Maier et al. 2020; Perumalsamy, Konidena, and Krothapalli 2023). While medical records and demographic factors can be used for mortality analysis, the inclusion of genetic factors can mean that risk prediction can be performed much earlier in an applicant’s life without necessarily the same amount of records as an older individual, emphasizing the risk of genetic discrimination (Karlsson Linnér and Koellinger 2022) based on the potential to have risk factors (e.g., a genetic risk for high blood pressure versus clinically-measured high blood pressure). Life insurance is not covered by GINA and life insurance companies can access medical records (which may include genetic test results) as part of an application. There is evidence of industry pressure and lobbying to shift the focus of bills to not protect genetic discrimination in life insurance (Rothstein 2018). Bills have been introduced in several states that would restrict the use of genetic information in underwriting, but these efforts largely failed: as of 2022, out of thirty-seven proposed bills across all states, three were introduced, eight were signed by the governor, demonstrating the geographical dependency of policy-based protection (Vermont Legislature 2019).
@@
Life insurance companies may be interested in using PGS to assess an individual’s risk for various diseases. In early 2024, U.S.-based life insurance company MassMutual and U.K.-based Genomics PLC announced a partnership, offering free genetic testing to MassMutual’s life insurance customers (MassMutual 2024). However, the press release stressed that MassMutual would not receive individual results and that current premiums/policies would be unaffected. This example particularly highlights the leaky interface between pillars, where the Right of Privacy and Right of Action are called into question. AI insurance underwriting algorithms already suffer from racial biases (Lee 2022). That PGS suffer from the portability problem (itself caused by systematic biases in training datasets) means that these biases could be perpetuated in genetics-informed life insurance underwriting. Additionally, given the relational aspect of genetic data, genetic-based underwriting could affect other biological relatives’ applications, harming not only relatives but any others with the same genetic markers that are associated with mortality risks. The vulnerabilities of (1) unsettled science and (2) rapid evolution and usages of genetic data can exacerbate the potential harms of using genetic data as part of underwriting in life insurance.
Recommendations
Recommendation 1: Redefining Genetic Data
Issue: Existing policy language does not adequately define genetic data or genetic information.
Recommendation: Given that genetic data possess unique characteristics distinct from other identifying information, we propose a two-tier definitional framework.
  1. 1.
@@
    Genetic data refers to any information relating to an individual’s genetic characteristics, including but not limited to DNA or RNA sequences, genomic variants, gene expression profiles, epigenetic markers, or genetic sequences from biological relatives. This data shall always be considered personally identifiable information (PII) as it pertains to unique biological attributes that can potentially be linked to a specific individual, their biological relatives, or identifiable group. For the purposes of this definition, any de-identified, pseudonymized, or anonymized genetic information shall be treated as genetic data, regardless of measures taken to mask individual identities, recognizing that the inherent characteristics of genetic data may enable re-identification through advanced technological or data cross-referencing methods.
  2. 2.
    Genetic information encompasses any downstream inferences, predictions, prior inferences, or derivations extracted from genetic data, including but not limited to health predispositions, ancestry determinations, trait predictions, and algorithmic outputs based on genetic data analysis.
While genetic information may not be immediately identifiable as traditional PII, we recommend that both genetic data and genetic information be treated as a distinct protected category for sensitive data requiring enhanced safeguards and the same level of protections as PII and protected health information (PHI). This approach draws from established legal precedent in biometric privacy law, specifically the distinction between “biometric identifiers” and “biometric information” established in Illinois’s Biometric Information Privacy Act (Illinois General Assembly 2008).
Recommendation 2: Extending Protections for Genetic Discrimination
@@
Issue: While GINA covers health insurance and employment, it does not protect life, long-term care, and disability insurance, as well as other domains (e.g., housing, education).
@@
Recommendation: We recommend that additional federal laws should be enacted to extend GINA’s coverage beyond employment and health insurance and for GINA to adopt and encompass the definitions of genetic data and genetic information provided in Recommendation 1. The most comprehensive state genetic anti-discrimination law is California’s CalGINA, which includes housing, mortgage brokerage, education, and more. However, despite these extensions, CalGINA does not cover life, disability, or long-term care insurance. We recommend that a comprehensive update to GINA should explicitly cover life, disability, long-term care, as well as education and any other opportunities for advancement to be protected against genetic discrimination. Legislation should be written in such a way that prohibits any barrier to opportunity, even those not anticipated at the time of writing. Additionally, it should explicitly bar genetic risk for complex traits known to have significant environmental influences (such as cardiovascular disease) from being considered a preexisting condition.
Recommendation 3: A Genetic Data Regulation Framework
Issue: Current protections do not comprehensively apply to the full genetic data ecosystem which leads to leaky protections (e.g., clinical tests being used in life insurance).